OpenStudy (anonymous):

\[\sum_{n=1}^{\infty} 2(\frac{ 9 }{ 10 })^{n}\] is there a sum to the series?

OpenStudy (anonymous):

there is in fact a sum to the series because r= 9/10 and IrI<1

OpenStudy (anonymous):

true story

OpenStudy (anonymous):

so it converges to 2?

OpenStudy (anonymous):

\[\sum_{0}^{\infty} r^{n}\] converges to \[\frac{ 1 }{ 1- r }\] if r < 1 we are dealing with \[2\sum_{0}^{\infty} (9/10)^{n}\ = 2*\frac{ 1 }{ 1 - (9/10) }\]

OpenStudy (anonymous):

so it would converge to 2/.10

OpenStudy (blurbendy):

isnt it 1 to infinity, not 0 to infinity?

OpenStudy (anonymous):

you're right blur

OpenStudy (anonymous):

oh it's when n=0

OpenStudy (anonymous):

and because n can get really big and anything to the ^1 = 1, the sum will be 2?

OpenStudy (anonymous):

my mistake; good observation blur. \[2\sum_{0}^{\infty - 1}(9/10)^{n-1} = 2\sum_{0}^{\infty}(9/10)^{n}*\frac{ 1 }{ (9/10) }\] \[= (20/9)\sum_{0}^{\infty}(9/10)^{n}\] rule applies

OpenStudy (anonymous):

So the sum of this series is 20?

OpenStudy (anonymous):

no it's: \[(20/9)*\frac{ 1 }{ 1 - (9/10) }\] . doens't look like it'll be an integer

OpenStudy (anonymous):

2.2 x 1.1 = 2.42?

OpenStudy (anonymous):

1/(1-9/10) doesn't = 1.1

OpenStudy (anonymous):

ok, thank you